Assume an individual’s preference for Fanta (Good x) and Fritters (Good y) are given by

the following utility function: U(x,y)=130= 3𝑥2 + 3𝑥𝑦 + 9𝑦

REQUIRED:

i) Derive the equation for the indifference curve for this utility function


1
Expert's answer
2021-09-14T09:30:33-0400

IC =>

"3x^2 + 3xy + 9y = 130 \\

3y(3+x) = 130 -3x^2 \\

y = \frac{130-3x^2}{3(3+x)}"